THE SPECTRAL STUDY OF THE OH (ν_{1}) + CH (ν_{3}) STRETCH COMBINATION BAND OF JET-COOLED METHANOL

\underline{SHUCHENG XU} and DAVID S. PERRY, Department of Chemistry, University of Akron, Akron, OH 44325.

The rotationally resolved overtone spectrum of the OH (ν_{1}) + CH (ν_{3}) stretch combination band of methanol between 6510 and 6550 cm^{−1} has been recorded at sub-Doppler resolution by continuous-wave cavity ringdwon spectroscopy (CW-CRDS). Of 572 recorded lines, 358 lines have been assigned by ground state combination differences, including the 21 subbands with J′ = 0-8 and K′ = 0-3. The perturbations of E K′ = -1, -2, and -3 levels have been observed and deperturbations have been carried out. The torsion-rotational constants in the upper state, obtained by fitting the observed spectrum to the Herbst Hamiltonian, are in a reasonable agreement with the predicted values based on the constants of the fundamental bands, ν_{1} and ν_{3}. This combination band has fewer perturbations than the fundamental ν_{1}, and the coupling matrix elements are smaller, 0.02 to 0.66 cm^{−1} as compared to 0.35 to 1.6 cm^{−1}.